OpenStudy (anonymous):

am i wrong anywhere? \[\frac{ 0 }{ 0 }=2\] \[\frac{ 100-100 }{ 100-100 }\] \[\frac{ 10^2-10^2 }{ 10(10-10) }\] \[\frac{ (10+10)(10-10) }{ 10(10-10) }\] \[\frac{ 10+10 }{ 10 }\] \[\frac{ 20 }{ 10 }\] \[\frac{ 0 }{ 0 } = 2\]

hartnn (hartnn):

u can't cancel (10-10) from numerator and denominator. because 10-10 = 0

OpenStudy (anonymous):

multiplication rule!! :O

OpenStudy (dape):

The fraction \(\frac{0}{0}\) is undefined, so setting it equal to 2 doesn't really makes much sense.

OpenStudy (anonymous):

why that doesnt work (10-10)/(10-10) = 0/0 you said 0/0=2 so (10-10)/(10-10)=0/0=2 so the resulting answer is not 0/0=2 but 0/0=4 and you keep on going 0/0=8,0/0=16 the only way you get 0/0=2 is if you assumed 0/0=1 but obviously 1 is not equal to 2 which is why you cannot divide 0 by 0